Circular No. 7595

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ab IN NGC 6130 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on unfiltered CCD images taken on Mar. 11.5 (mag ∼ 17.5) and 12.5 UT (mag ∼ 17.4) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001ab is located at α = 16^{h}19^{m}34^{s}.83, δ = +57^{o}37′05′′.5 (equinox 2000.0), which is 10′′.9 east and 11′′.4 north of the nucleus of NGC 6130. A KAIT image taken on 2000 Aug. 15.3 showed nothing at this position (limiting mag ∼ 19.0).

V4643 SAGITTARII A. Nakamura, Kuma, Ehime, Japan, reports the following magnitudes

for V4643 Sgr by A. Tago (Tsuyama, Okayama), which he derived (using GSC comparison stars) from his T-Max 400 film exposures obtained with a 105-mm f/4 camera lens: Feb. 19.863 UT, [10.0; 22.860, [9.8; 26.870, 9.5.

GRS 1758–258 D. M. Smith, Space Sciences Laboratory, University of California at

Berkeley; C. Markwardt, Goddard Space Flight Center (GSFC); W. A. Heindl, Center for Astrophysics and Space Science, University of California at San Diego; and J. H. Swank, GSFC, using data from the RXTE PCA, report: “The black-hole candidate and microquasar GRS 1758–258 has entered an extremely dim state. The power law (usually photon index −2), which normally dominates the spectrum, has dropped by an order of magnitude in luminosity (from 10^{−9} to 10^{−10} erg cm^{−2} s^{−1}, 2–20 keV) and softened to index −3, leaving the spectrum dominated by the faint thermal component that is often present (kT = 0.4 keV). The change occurred abruptly: the power law was as bright as usual on Feb. 21, and on Feb. 27, Mar. 2, and 6 was in the dim state, with little apparent change among the latter three observations. We identify this with the ‘off’ state last reported by SIGMA in 1991–1992 (Gilfanov et al. 1993, Ap.J. 418, 844). The remaining soft emission may be the extended emission reported by Chandra from this source (Heindl and Smith 2000, Bull. A.A.S. 32, 1603). Observations are strongly encouraged at all wavelengths not excessively absorbed near the Galactic center, both to characterize this x-ray state and to search for changes in the optical candidates for this source’s companion, now that the accretion disk may have turned off.”

2001 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green